Consider the UNILUS economy in the long-run described by the following equations:

Y = C + I + G, Y = 5,000, G = 1,000, T = 1,000, C = 250 + 0.75(Y − T), I = 1,000 − 50r;

A. Compute private saving, public saving, and national saving.

(5 Marks)

B. Find the equilibrium interest rate. (5 Marks)

C. Now suppose that G rises to 1,250. Compute private saving, public saving, and national savings. (5 Marks)

D. Find the new equilibrium interest rate. (5 Marks)


1
Expert's answer
2021-09-07T08:00:02-0400

"Private savings (S) = Y \u2013 T \u2013 C"

"= 5000-1000 -( 250 + 0.75(5000-1000)""= 4000-3250"

"= 750"


"Public savings = (T \u2212 G)"

"1000- 1000"

"= 0"


"National Savings = private savings + Public savings""= 750+0= 750"